### v1

# Assessing realistic binding energies of some essential interstellar radicals with amorphous solid water:

11email: milansil93@gmail.com 22institutetext: Univ Rennes, CNRS, IPR (Institut de Physique de Rennes) - UMR 6251, F-35000 Rennes, France 33institutetext: Department of Chemical Sciences, Indian Institute of Science Education and Research Kolkata, Mohanpur 741246, West Bengal, India
33email: aroy.astro@gmail.com 44institutetext: Rosseland Centre for Solar Physics, University of Oslo, PO Box 1029 Blindern, 0315 Oslo, Norway 55institutetext: Institute of Theoretical Astrophysics, University of Oslo, PO Box 1029 Blindern, 0315 Oslo, Norway 66institutetext: Institute of Astronomy Space and Earth Sciences, P 177, CIT Road, Scheme 7m, Kolkata 700054, West Bengal, India
66email: ankan.das@gmail.com 77institutetext: Department of Chemistry, Graduate School of Science and Engineering, Tokyo Metropolitan University, 1-1 Minami-Osawa, Hachioji, Tokyo 192-0397, Japan 88institutetext: Institute of Science and Technology, Niigata University, Ikarashi-nihoncho 8050, Nishi-ku, Niigata 950-2181, Japan 99institutetext: National Astronomical Observatory of Japan, Osawa 2-21-1, Mitaka, Tokyo 181-8588, Japan 1010institutetext: Department of Astronomy, Graduate School of Science, University of Tokyo, Tokyo 113-0033, Japan 1111institutetext: Universidad Autónoma de Chile, Facultad de Ingeniería, Núcleo de Astroquímica & Astrofísica, Av. Pedro de Valdivia 425, Providencia, Santiago, Chile 1212institutetext: Max-Planck-Institute for extraterrestrial Physics, P.O. Box 1312 85741 Garching, Germany
Assessing realistic binding energies of some essential interstellar radicals with amorphous solid water:
In the absence of laboratory data, state-of-the-art quantum chemical approaches can provide estimates on the binding energy (BE) of interstellar species with grains. Without BE values, contemporary astrochemical models are compelled to utilize wild guesses, often delivering misleading information. Here, we employ a fully quantum chemical approach to estimate the BE of 7 diatomic radicals – CH, NH, OH, SH, CN, NS, and NO, playing a crucial role in shaping the interstellar chemical composition, using a suitable amorphous solid water model as a substrate since water is the principal constituent of interstellar ice in dense and shielded regions. While the BEs are compatible with physisorption, the binding of CH in some sites shows chemisorption, in which a chemical bond to an oxygen atom of a water molecule is formed. While no structural change can be observed for the CN radical, it is believed that the formation of a hemibonded system between the outer layer of the water cluster and the radical is the reason for the unusually large BE in one of the binding sites considered in our study. A significantly lower BE for NO, consistent with recent calculations, is obtained, which is useful in explaining the recently observed HONO/NH2OH and HONO/HNO ratio in the low-mass hot corino, IRAS 16293-2422B, with chemical models.
Key Words.:
Astrochemistry – Interstellar medium (ISM) – ISM: molecules – ISM: structure – ISM: dust – ISM: abundances – evolution – Molecular processes1 Introduction
Currently, more than 300 molecules have been detected in the interstellar medium or circumstellar shells111https://cdms.astro.uni-koeln.de/classic/molecules (McGuire, 2022). Interstellar grains act as catalysts in shaping the chemical composition of the interstellar clouds (Hasegawa et al., 1992; Das et al., 2008a, 2010, 2016, 2019; Das & Chakrabarti, 2011; Srivastav et al., 2022; Sil et al., 2018, 2021; Ghosh et al., 2022; Mondal et al., 2021). is the most dominant ice component in dense and cold regions of interstellar clouds (Gibb et al., 2004; Boogert et al., 2015). It can account for the % of the ice along most lines of sight (Whittet, 2003). In dense and shielded regions, multiple layers of ice would form either by the direct adsorption of gas phase species or by chemical processing on the grain surface. The proper structure of interstellar ice remains unclear, although it is generally agreed that it is amorphous upon formation (Hama & Watanabe, 2013).
Estimating binding energies (hereafter BEs) of interstellar species is essential to interpret observations and reconstruct our astrochemical history. Several recent studies attempted to provide theoretical BEs of interstellar species. Sil et al. (2017) calculated the BE of most abundant H and on different types of grain surfaces and found that the BE of is always higher than that of H; Wakelam et al. (2017) reported the physisorbed BE of more than interstellar species, including radicals, using a single molecule (monomer) as an approximate amorphous solid water (ASW) ice surface based on quantum chemical calculations. They compared the calculated BEs with experimental values for some molecules. They concluded that the calculated values of the physisorbed BE considering a single molecule are proportional to the experimental values. Das et al. (2018) attempted to improve on this using different sizes of clusters containing up to 6 molecules to calculate the BEs of species, including radicals of astrochemical interest and noted that with increasing cluster size, the calculated BE values start to converge toward the experimentally obtained values. Shimonishi et al. (2018) introduced molecular dynamic (MD) simulations for the first time to properly describe the ASW cluster of 20 molecules and their interactions with atomic C, N, and O using quantum chemistry. Ferrero et al. (2020) employed density functional theory (DFT) using B3LYP-D3(BJ) (for closed-shell molecules) and M06-2X (for radical species) functionals and Ahlrichs’ triple-zeta quality VTZ basis set, supplemented with a double set of polarization functions (A-VTZ*) to calculate the BEs of molecules and radicals (consisting of up to binding sites per each), considering both crystalline and amorphous periodic ice (consisting of ) models. They found good agreement with the previously mentioned studies. In a similar vein, Perrero et al. (2022) and Martínez-Bachs et al. (2024) investigated the BEs of various S- and N-bearing molecules and radicals adsorbed on similar water ice models using similar DFT and higher-level composite methodological approaches. Tinacci et al. (2022) developed a computational framework to simulate the formation of interstellar ice mantles through H2O accretion. They used the ONIOM approach to compute the BE distribution of NH3 considering the ASW ice model composed of 200 H2O molecules. Their results highlighted the complexity of adsorption sites and the potential for multiple BE regimes. In a later study, Tinacci et al. (2023) calculated the distribution of the H2O BE on the same ASW grain model and integrated it into a protoplanetary disk model. They assessed its effect on the position of the water snow line and estimated the water content of potential planetesimals at different distances from the disk center. Bariosco et al. (2024) expanded on this by investigating the distribution of the H2S BE on the same ASW ice surfaces, emphasizing the importance of considering a multitude of binding sites. Duflot et al. (2021) applied the ONIOM high-quantum-mechanic (QM) part and a low-QM part hybrid method to calculate the BEs for a series of atoms, small molecules, and radicals on both crystalline and amorphous ice represented by a cluster of about molecules obtained via classical MD and electronic structure methods. Sameera et al. (2021) used the ONIOM QM/MM method for binding sites of the radical BEs calculation on both crystalline and amorphous ice clusters containing molecules ( in the QM and in the MM region). Enrique-Romero et al. (2022) computed the BEs of radicals using DFT calculations on two ice water cluster models made of and molecules. Bovolenta et al. (2022) computed BE distributions of interstellar molecules and radicals on an amorphized set of water clusters of molecules each, containing unique binding sites. Piacentino & Öberg (2022) published the BEs of P-bearing species on molecule clusters using DFT (M06-2X/aug-cc-pVDZ level of theory). Hendrix et al. (2023) reported computational searches for optimal structures and benchmark binding and condensation energies for sets of neutral, radical, cationic, and anionic molecules of astrochemical relevance with clusters of N molecules. They demonstrated excellent agreement (with discrepancies less than K) between the DFT (B97X-V/def2-svpd and B97X-V/def2-qzvppd level of theories) and coupled-cluster (CCSD(T)/aug-cc-pVTZ level of theory) BEs, which validates the DFT protocols. This could make the DFT level of theory the primary method for studying larger clusters.
The aforementioned studies prevalently focus on the closed-shell species, where all the electrons are paired. In contrast, open-shell species (like radicals) pose technical challenges due to convergence issues in self-consistent field (SCF) procedures. The energy surface of radicals is more complex, with multiple local minima, making it harder to find the global minimum and the unpaired electrons result in a high electron correlation effect, requiring a more sophisticated treatment. Experimentally determining the desorption data of unstable radicals is more challenging due to difficulties in their preparation and deposition, and their tendency to react before desorbing. The chemical energy released upon reaction can also trigger a desorption event (Minissale et al., 2022). Therefore, focusing on stable species is more common due to the difficulties in handling molecular radicals experimentally.
The present work provides the BEs on various amorphous water ices for some open-shell neutral radicals that can be directly implemented into astrochemical models. In the past, there was a prevailing belief that radicals only diffused at relatively high temperatures (¿30 K; Garrod et al., 2008; Das et al., 2015). However, contemporary research suggests that complex organic molecules (COMs) can be generated through non-diffusive chemistry (Shingledecker et al., 2019; Jin & Garrod, 2020; Ioppolo et al., 2021). Since water is the major constituent of a grain mantle in molecular clouds (Keane et al., 2001; Das et al., 2010; Das & Chakrabarti, 2011; Das et al., 2016), we perform BE computations of the radicals CH, NH, OH, SH, CN, NS, and NO considering 9 different ASW clusters constructed with 20 molecules (). A detailed comparison is also made between the calculated BEs and previously available experimental or theoretically obtained BE values.
2 Computational details and methodology
The BE is usually seen as a local property arising from the electronic interaction between the substrate (grain surface or adsorbent) and the species deposited on it (adsorbate). For a bounded adsorbate, the BE is a positive quantity. It is defined as the difference in electronic optimized energy between the separated radical () and substrate () and the adsorbed radical on the substrate (), i.e.,
For our BE investigation, we consider nine different structures of ASW clusters (see LABEL:tab:opt_water) as the adsorbents. Each ASW ice model was created by cutting out a block from a large unit cell of crystalline ice. They are heated to 300 K using molecular dynamics (MD) annealing calculations based on classical force fields. To prevent water molecules from escaping the cluster, 20 water molecules using the TIP3P model are selected, and droplet simulations are run with a spherical restriction of 20 . Following 100 ps simulations at 300 K, the nine distinct structures are rapidly cooled to a low temperature of 10 K (Shimonishi et al., 2018). These ASW structures are a good compromise between accuracy and computational time/cost for several radicals. Also, ice in astrophysical environments, such as icy satellites, comets, planetary rings, interstellar grains, etc., are mainly amorphous, and the main component of the ice mantle is (Boogert et al., 2015). Moreover, some recent studies (Germain et al., 2022; Tinacci et al., 2022) suggest that a large number (at least a few hundred) of molecules in the ASW cluster could mimic the interstellar ice mantle. However, these types of ASW are too large to handle the complete quantum chemical treatment (geometric optimization and harmonic frequency evaluation). Considering a large number of binding sites, cluster would be a reasonable choice to study the BE. The smaller size of ASW clusters results in the formation of ?under-coordinated? water molecules at their surface, thus pushing up their affinity toward adsorbates and higher BE values. The larger ASW clusters, on the contrary, tend to maximize the number of internal H-bonds. This gives them a more spherical shape and thus has fewer available strong adsorption sites for the adsorbates (Martínez-Bachs et al., 2024) due to a reduction in edges and/or cavities in the model.
All the quantum chemical calculations are performed using the Gaussian 09 and Gaussian 16 suite of programs (Frisch et al., 2009, 2013, 2016, 2019). The considered nine different structures of clusters are fully optimized based on DFT calculations using B97X-D functional (Chai & Head-Gordon, 2008) approach, in which the van der Waals (vdW) interaction (dispersion correction) is empirically incorporated. The choice of B97X-D was motivated by its demonstrated accuracy in describing systems with unpaired electrons, such as radicals. The range-separated hybrid nature of this functional enables a balanced treatment of both electrostatic and dispersion interactions, which are crucial for accurately modeling radical-water complexes. While benchmarking against higher levels of theory is beyond the scope of this work, we acknowledge its significance and intend to address this in future studies. A valence triple- plus polarization and diffuse functions, namely 6-311+G(d,p) basis set, is adopted for the geometry optimization and the BE evaluation. A fully optimized electronic ground state structure, i.e., the local minimum, is verified to be a stationary point (having non-negative frequency) by harmonic vibrational frequency analysis. All the BEs are evaluated with the harmonic vibrational zero-point energy (ZPE) corrections to consider the molecular vibration at low temperatures. Unless otherwise stated, all the BEs reported or compared in this work are the ZPE-corrected BEs. The basis set superposition error (BSSE) is taken care of by the counterpoise correction (CPC) method (Boys & Bernardi, 1970) to get an idea of the BSSE dependency of the system (see 3.2).
We use clusters to simulate the interstellar ASW surface and investigate the influence of H-bond cooperativity on radical BEs as well as the effect of the BE on binding site variation (Bovolenta et al., 2020; Ferrero et al., 2020; Duflot et al., 2021; Bovolenta et al., 2022). The accessible radical binding sites are chosen based on the available dangling-H (-H) positions on ASW cluster surfaces. The BE is controlled by electrostatic attractions, orbital interactions, and Pauli repulsion. As a result, each radical species can have a range of possible BEs (Minissale et al., 2022). As per the information provided in LABEL:tab:opt_water, it is seen that a total of 75 -H sites are available; these sites are distributed among ASW cluster 1 to 9, having 7, 7, 9, 10, 8, 9, 8, 9, and 8 -H sites, respectively. Using a weak H-like bond, geometry optimizations of the complex structures are performed after manually connecting the radical with the -H. During this weak bond connection, we always prefer the largest differences in electronegativity values (see Table 1) between two binding elements of the atom in the radical and -H atom in the cluster. The electronegativity of each atom contributes to the formation of molecular dipoles, which can be strategically utilized to construct the initial adsorption geometries manually. By following the principle of electrostatic complementarity, we always prefer to connect the -H atom of the cluster with the atom in the radical that has the higher electronegativity. This approach leverages the natural tendency of atoms with higher electronegativity to attract electrons or electron density toward themselves, resulting in stronger dipole interactions and more stable adsorption configurations. The electronic ground state spin multiplicity of each of the 7 radicals was checked and noted in Table 2. All the considered radicals are doublets (with one unpaired electron) except NH, which is in the triplet (with two unpaired electrons of the same spin) state. The coordinates of all the representative structures are publicly available in a GitHub repository222https://github.com/milansil/ASW_radicals_data.git.
Species | Electronegativity |
H | 2.1 |
C | 2.5 |
S | 2.5 |
N | 3.0 |
O | 3.5 |
It is to be noted that the optimized energy and the calculated BEs obtained from the Gaussian program are not always accurate due to stationary point convergence in frequency domain calculations. This is because the root-mean-square (RMS) and maximum force are calculated analytically in the self-consistent field (SCF) and frequency domain. However, SCF numerically solves the maximum and RMS displacement; on the contrary, the post-SCF (frequency) solves both of them analytically. Hence, some deviation of K can be observed. Due to time constraints and high computational cost, the values are taken even when only the maximum and RMS forces are being converged in the frequency domain, and the frequencies are all positive.
3 Computational results and discussions
3.1 Binding energy
Firstly, we calculate the optimization energy of 9 ASW clusters shown in LABEL:tab:opt_water along with their optimized structures. It is seen that the ASW(1) cluster is, on average, K more stable than the other 8 ASW clusters. We then calculate the BE of OH radical at the -H binding sites of 9 ASW clusters. All the calculated (+) BSSE BE values are plotted in Fig. 1 and tabulated in LABEL:tab:BE_OH.
Instead of taking the average BE over all the binding sites, we consider the average over the highest -H binding sites in each cluster. This averaging would be more appropriate for estimating the BE at a low temperature (interstellar condition). As a result, we obtain average BE values of K and K for the OH radical with and without considering BSSE corrections, respectively. Depending upon the binding sites, the BEs are rather broadly distributed (see LABEL:tab:BE_OH). To verify the distribution of the BSSE-corrected BE values over the highest -H binding sites in each cluster, we organize them in a distribution plot with a bin width of 500 K shown in Fig. 2. The distribution is well reproduced by an overlay of Gaussian Probability distribution function:
where is the mean and is the standard deviation. The fitted parameters for the BE values distribution are the mean () K and standard deviation () K. It implies that the Boltzmann distribution (uniform distribution to the highest BE site at K) holds for each local region, which is further supported by the fitting using the Boltzmann model equation (Fig. 7). On the contrary, the BE distribution considering all the 75 sites (BSSE-corrected and uncorrected) does not mimic a Boltzmann distribution (non-Gaussian type), as depicted in Fig. 8.
To minimize the computational burden, we perform the calculation of the BE for the remaining 6 radicals (i.e., CH, NH, SH, CN, NS, and NO) by considering only the ASW(1) cluster. This particular ASW cluster is the most stable (lowest energy) among the 9 ASW clusters we have mentioned earlier (see LABEL:tab:opt_water). Then, we select the highest physisorbed BE values among 7 binding sites available in ASW(1) (see Table 7). It is interesting to note that except for NO, all the other 6 radicals (i.e., CH, NH, OH, SH, CN, and NS) give the highest BEs (considering both physisorption and chemisorption) with the fifth -H position of ASW(1) cluster. Since we are only considering ASW(1) cluster to estimate the BE of the 6 radicals except for OH, we want to examine the effect on the average BE when considering all 9 ASW clusters. For OH, we have both the average BE considering all 9 ASW clusters and the maximum BE from ASW(1) cluster. Hence, we calculate a ratio of 1.177, indicating around deviation, obtained from OH radical calculation by dividing the BSSE-corrected average BE (4287 K) by the BSSE-corrected highest BE (3643 K) of ASW(1) cluster (see LABEL:tab:BE_OH). Consequently, we apply this scaling factor to the other 6 radicals to determine the effect on the average BE over all the 9 ASW clusters, as we only estimated the BEs for these species using the ASW(1) cluster. As mentioned earlier, the standard deviation ( 566 K) for OH is calculated using the Gaussian distribution over all the 9 highest BEs obtained for the 9 ASW clusters, and thus 566 K can be treated as an error bar for the BE of OH. To get an idea of this error bar for the other 6 radicals, a standard deviation ( 786 K) for OH is calculated by taking the 7 binding sites of the ASW(1) cluster. Finally, a scaling factor of is obtained by taking their ratio. For the other 6 radicals, the standard deviation is obtained similarly as for OH radical considering the 7 binding sites of the ASW(1) cluster and then multiplied with the scaling factor of to get an estimate of the standard deviation considering the 9 ASW clusters (see Fig. 11). It is to be noted that the calculation does not take into account the significantly higher BE values for CH and CN, as these are considered special cases. The final BSSE-corrected scaled BE values, along with the error bars, are noted in the last column of Table 2.
Sl. | Radicals | Formula | Ground state | BE in Kelvin | ||||
No. | spin multiplicity | PENTEADOa | DASb | KIDA∗ database | Experimental | This work | ||
1. | Methylidyne radical | CH | Doublet | — | — | |||
2. | Imidogen | NH | Triplet | 1947 | — | |||
3. | Hydroxyl radical | OH | Doublet | 3183 | ||||
4. | Mercapto radical | SH / HS | Doublet | 2221 | — | |||
5. | Cyano radical | CN | Doublet | 1736 | — | |||
6. | Nitrogen monosulfide | NS | Doublet | 2774 | — | |||
7. | Nitric oxide | NO | Doublet | — |
3.2 BSSE correction
The basis set superposition error (BSSE) is an intuitively highly plausible concept. The BSSE correction using the CPC method accounts for the superposition error caused by the sloppy treatment of the basis set for each monomer of a big cluster, as the intermolecular distance varies. Eventually, even if this inconsistency could be eradicated entirely, errors related to the incompleteness of the basis set, known as ?basis set completeness error? (BSCE), would be preserved. Hypothetically, the BSSE and the BSCE would be zero in the complete basis set (CBS) limit. The BSSE and BSCE often have opposite signs (Dunning, 2000). For small and medium basis sets around the vdW minimum, the (negative) BSCE cancels (part of) the (positive) BSSE (Sheng et al., 2011). Ignoring one basis set fault while accounting for the other can produce imbalanced results. According to Sheng et al. (2011), diffuse basis functions can help reduce the BSSE. However, they also pointed out that diffuse basis functions can reduce the BSCE. Therefore, the partial cancellation of these errors still applies, and the extensive use of diffuse basis functions does not lead to a different conclusion. Jensen (2024) recently pointed out that the most significant source of error for smaller basis sets is BSCE. Furthermore, he noted that these errors are dominant in dispersion-type interactions and that including CPC may, in such cases, lead to a deterioration of the results.
The recent studies by Perrero et al. (2022); Bovolenta et al. (2022); Tinacci et al. (2023) suggest that the BE values can be erroneous for bigger systems without the BSSE correction. Considering both the contradictory statements and also to reduce the computational cost and make the calculations possible, we have used only one small basis set, 6-311+G(d,p), and thus cannot extrapolate toward the CBS limit, hence an estimation has been made to determine the influence of BSSE on the highest physisorption BEs of the radicals.
Sl | Radicals | B97X-D / 6-311+G(d,p) | ||
No. | BSSE | BSSE | ||
1. | CH | 1892 | 1737 | 1.0892 |
2. | NH | 2083 | 1898 | 1.0975 |
3. | OH | 3990 | 3643 | 1.0952 |
4. | SH / HS | 2416 | 2109 | 1.1456 |
5. | CN | 1192 | 1086 | 1.0976 |
6. | NS | 2606 | 2225 | 1.1712 |
7. | NO | 931 | 598 | 1.5569 |
In Table 3, we compare the results obtained by considering BSSE correction using the CPC method and BSSE-uncorrected BE for the -H binding sites with the highest BE found for all the 7 radicals with ASW(1) cluster. It is to be noted that for the CH radical, the highest BEs are due to chemisorption process (see Table 7). Chemisorption refers to forming a chemical bond between two atoms of the radical and water cluster. The structural geometries (see Fig. 9) suggest that for CH radical, sites 5 and 7 in ASW(1) cluster, the cluster geometries get broken, and consequently, CH2OH and CHOH+H formation happens respectively. It is pretty similar to the case found for the C and CH in Wakelam et al. (2017); Shimonishi et al. (2018); Molpeceres et al. (2021). We consider the highest BE value among only the physisorbed energy values to keep calculation consistency. The result is less evident for CN radical (exceptionally large BSSE-corrected BE of 6317 K with site 5 in ASW(1) cluster) as no structural change can be seen. It is thought that CN forms a non-classical 2c-3e bond (hemibond) between the unpaired electron on the CN carbon atom and the lone pair electrons of the oxygen atom of the water surface in place of forming H-bonds with the surface (Rimola et al., 2018; Martínez-Bachs et al., 2024) and is most probably the reason behind the significantly large BE value. For the case of CN, we consider 1086 K as the BSSE-corrected highest BE value for consistency. To better understand the dependency of BSSE on the BE, the BSSE-corrected and uncorrected BEs are plotted side by side and compared along with the ratio by which the BSSE-corrected BEs are lower than the uncorrected one for all the 7 radicals in Fig. 3. From Table 3, we find that the average ratio , considering the -H binding sites with the highest BE found for all 7 radicals with ASW(1) cluster. A quantitative study on the BSSE dependency of the OH radical considering all the binding sites in ASW(1) cluster is also performed and tabulated in Table 8. The BSSE-corrected and uncorrected values and uncorrected to corrected ratio are plotted in Fig. 10. The uncorrected to corrected average ratio for OH is . It is seen that the BE values with BSSE adjustments are consistently lower than those without BSSE corrections for all the cases. Unless otherwise stated, all the BEs reported or compared afterward are the BSSE-corrected BEs.
3.3 Comparison of the BEs with previous studies
The experimental investigation of radical binding properties is complex since, during the heating ramp of conventional temperature programmed
desorption (TPD), the radical has an enormous chance to diffuse and react, making the experimental evaluation of the desorption parameters almost impossible.
The obtained BE values depend strongly on the chemical composition and morphology of the substrate, as well as the monolayer or multilayer regime used in the experiment (He et al., 2016).
To the best of our knowledge, no experimental studies are available to provide the BEs of radicals considered in this study, except for a handful of experimental determinations for the BE of OH on amorphous silicate surfaces.
Therefore, we do not have many constraints to benchmark our computed values.
Moreover, comparing the theoretical BEs with the experiment is not straightforward.
We must, therefore, rely on the other theoretical existing reference values.
In Table 2, we compare our calculated BE with other available experimental, computational, and public database values.
To better visualize the comparison, we plot all the tabulated data on a common diagram in Fig. 4.
To use a coherent set of homogeneous data for comparison, we chose the recommended BE values by Das et al. (2018) except for CH and Penteado et al. (2017) and frequently used the KIDA database for all the considered radicals.
Penteado et al. (2017) presented a list of the recommended BEs, which includes data from previous studies and an uncertainty range for each value.
The uncertainty assigned is set to half the BE for species with less than 1000 K (CH and NH) and 500 K for all other cases (SH, CN, NS, and NO).
For all the radicals considered here except OH, the provided average BEs are based on the work of Hasegawa & Herbst (1993); Aikawa et al. (1996).
OH: The binding properties of the OH radical are difficult to study experimentally due to its tendency to diffuse and react, leading to the formation of ASW on the bare surface during conventional TPD heating ramps. Only a couple of existing experimental studies reported the BE for hydroxyl radical on bare amorphous silicates as 4600 K (Dulieu et al., 2013) and K (He & Vidali, 2014). Penteado et al. (2017) later recommended BE = K, based on the distribution determined by He & Vidali (2014), taking the average with uncertainty to cover this range.
However, as far as we know, no experimental study considering ASW surface has been published to date.
Several other existing reference values are available for OH.
The OSU gas-phase database provides 2850 K, assuming half the value of (Garrod & Herbst, 2006).
A systematic empirical approach by Wakelam et al. (2017) provided K considering ASW model.
Das et al. (2018) computed by quantum calculations 3183 K considering ASW model.
Ferrero et al. (2020) determined a range of values K considering the amorphous 2D slab ice model, generated from the structures of Shimonishi et al. (2018).
There are some estimated average values ranging between 4300 K (considering both crystalline and ASW substrates; Miyazaki et al., 2020) and 4990 K (between 2321 K and 7770 K on top of a crystalline water ice; Sameera et al., 2017).
Duflot et al. (2021) found average best estimate of K between minimum 2936 K and maximum 8924 K considering amorphous ice model.
Enrique-Romero et al. (2022) calculated 2911 K and 5376 K on the and ASW ice model.
Finally, after examining the available studies, Minissale et al. (2022) recommended a value of 5698 K.
Very recently, Hendrix et al. (2023) computed the BEs of OH to monomer, dimer, trimer, and tetramer as 1922 K, 3900 K, 4252 K, and 3674 K, respectively using CCSD(T)/aug-cc-pVTZ level of theory and 1907 K, 3905 K, 4433 K, and 3840 K, respectively using DFT-B97X-V/def2-qzvppd level of theory.
They noticed that the H-bond distances between the O atom of OH and the H atom of shorten notably upon adding a molecule to the cluster.
It is seen that from Fig. 4, our calculated average BE of K is well within the range of existing experimental, theoretical, and recommended values and close to the central value.
CH: As per our knowledge, no computational or experimental BE value exists for methylidyne (also called carbyne) radical to date.
Wakelam et al. (2017) attempted to consider CH in their work but noticed the formation of strong (partly) covalent bonding with a single H2O molecule. However, they did not provide any chemisorption value. We calculate its BE considering 7 -H sites of ASW(1) cluster (see Table 7).
[…]
A recent study by Nguyen & Peeters (2021) on the reaction rate study of CH with H2O vapor suggests that in the reaction pathway, the first step is the barrier-less association
leading to a pre-reactive complex (PRC), which has a BE of kcal mol-1 ( K). This PRC (HCδ+–δ-OH2) is formed by
a polar interaction of a lone pair orbital of the O atom and an empty p-orbital of the C atom. Subsequently, the PRC can isomerize to CH2OH by insertion of CH into an O–H bond of H2O.
A similar chemisorption behavior of C-atom with H2O was recognized by previous studies (Wakelam et al., 2017; Shimonishi et al., 2018; Molpeceres et al., 2021).
The large dipole moment (1.46 D; Phelps & Dalby, 1966) could be the reason behind the strong long-range interaction and thus showing reactivity toward neutral molecules.
As seen from Table 2 and Fig. 4, our proposed average BE of K is larger by comparing the available estimated (based on educated guesses) range of K provided in Penteado et al. (2017), as well as 925 K provided by the OSU gas-phase database.
To clarify, our average BE is calculated without taking into account the chemisorbed BE values (see Table 7).
NH: We report the average BE for imidogen or nitrene as K, which is in good agreement with the OSU database value of 2378 K and within the range of values of K proposed by Wakelam et al. (2017).
Our calculated value is well above the recommended value of K (Penteado et al., 2017) and larger than the computed value of 1947 K by Das et al. (2018).
Martínez-Bachs et al. (2020) reported the BEs of 4222 K considering crystalline water-ice surface.
Our value agrees well with the best-estimated average value of K considering the amorphous ice model (Duflot et al., 2021).
Enrique-Romero et al. (2022) have recently calculated the BEs of NH as 1560 K and 3909 K on the and ASW ice models respectively considering a single binding site which is indeed a simplistic assumption given that different surface binding sites are available and, accordingly, a distribution of BEs exist.
Enrique-Romero et al. (2022) also noted (from the private communication with B. Martinez-Bachs) the range of BEs of K considering ASW ice surfaces, centered at around 2410 K.
However, the very recent calculations by Martínez-Bachs et al. (2024) show a range of BEs on the ASW model varying from a minimum of 1167 K to a maximum of 4692 K, with an average of 2464 K, which also gives a strong agreement with our value.
SH/HS: We report the average BE for mercapto or sulfanyl radical as K, which is within the range of values of K proposed by Wakelam et al. (2017).
However, the other values 1450 K (OSU database), K (Penteado et al., 2017), 2221 K (Das et al., 2018), and K (Perrero et al., 2022) are found somewhat smaller than our computed average value.
CN: Our computed average BE value for the cyano radical ( K) is consistent with the existing values of 1600 K (OSU database), K (Penteado et al., 2017), and 1736 K (Das et al., 2018), except the larger value K proposed by Wakelam et al. (2017).
However, an exceptionally strong BE of 6317 K is found for site 5 in ASW(1) cluster. This value is not included in the average calculation. Eventually, this high BE value coincides with the range of BEs proposed by Martínez-Bachs et al. (2024), with the values ranging from 2605 K to 7792 K and an average of 4856 K. The hemibond nature is most probably the acceptable reason behind this inconsistency of the BE values.
NS: Unlike the case of the CH radical, our calculated average BE for the nitrogen monosulfide radical K is larger by comparing the available estimated range of K (Penteado et al., 2017), as well as 1900 K provided by the OSU gas-phase database.
Das et al. (2018) calculated value of 2774 K is somewhat close to our calculated average value.
Perrero et al. (2022) calculated the average BE of 2026 K between minimum 1234 K and maximum 2955 K considering an amorphous ice model.
NO: Our computed average BE value of K for nitric oxide radical is well within the range K proposed by Penteado et al. (2017), and the calculated value ( K) on different ASW ice model by Das et al. (2018). The KIDA database suggests a higher value of K as proposed by Wakelam et al. (2017).
Very recently, Hendrix et al. (2023) computed the BEs of NO to monomer, dimer, trimer, and tetramer as 367 K, 1147 K, 438 K, and 1354 K, respectively, using CCSD(T)/aug-cc-pVTZ level of theory and 252 K, 1022 K, 649 K, and 1087 K, respectively, using DFT-B97X-V/def2-qzvppd level of theory. Similar to NH, the results of Martínez-Bachs et al. (2024) give an energy range of 291 K to 965 K, with an average of 712 K, thus strongly agreeing with our results. The weak BE of NO on the surface is primarily governed by dispersive interactions. This is due to the almost negligible dipole moment of NO, which limits the contribution of electrostatic interactions to the overall binding.
It is important to note that the small fluctuations observed in the calculated BEs compared to literature values can be attributed to several factors, including the specific computational method employed, the adopted ice model, and the particular binding sites investigated in this study.
Open-shell species pose significant computational challenges due to multireference characteristics and the need to account for static and dynamic correlation effects. Accurately capturing the true energy of these species requires a sensitive and careful approach, often involving multireference wave function theories (e.g., coupled cluster method) and perturbation methods (e.g., Møller-Plesset method). Although these techniques can provide highly accurate results, they come with a substantial computational cost. The combination of the B97X-D functional and the 6-311+G(d,p) basis set offers a strong solution to achieve a practical balance between accuracy and efficiency. The B97X-D functional is crucial for handling the complex electronic structures inherent to radicals and is a versatile tool for studying a wide range of systems. The 6-311+G(d,p) basis set, with its valence triple- quality and added polarization functions, further enhances the accuracy by providing a flexible and detailed description of electron density, especially where electron distribution is uneven due to unpaired electrons. This combination is not necessarily the best or most accurate option available. Still, it is a well-considered choice that delivers a high degree of accuracy while keeping computational costs relatively low. Furthermore, we believe our methodology with careful structural optimization and analysis is robust enough to be applied to other radical species and, more generally, to species with challenging electronic structures.
4 Astrophysical Implications
It is crucial to comprehend the impact of the newly computed BEs on interstellar ice chemistry, as some of them differ from the values used in the past few decades. We outline the implications of these computed values below.
4.1 Chemisorption
Some of our binding sites exhibit chemisorption for CH (a chemical bond is formed with one oxygen atom of a water molecule of the cluster). Figure 9 depicts the formation of CH2OH and CHOH2 by chemisorption. There is a possibility that the chemisorbed CH2OH could produce CH3OH by the Eley-Rideal mechanism by capturing an H atom from the gas phase. Alternatively, a chemisorbed H atom could react with CH2OH to form methanol. Recent experiments by Grieco et al. (2024) have suggested the formation of H2O at a high temperature (approximately 85 K) by chemisorbed H and O. Additionally, Grieco et al. (2023) reported the formation of H2 up to a temperature as high as 250 K, suggesting the feasibility of the hydrogenation reaction at higher temperatures than previously thought. Contrary to previous understanding of methanol formation through the surface hydrogenation of CO, Molpeceres et al. (2021) and Tsuge et al. (2023) suggest that before CO freezeout, atomic carbon deposited on ASW could form H2CO. Tsuge et al. (2023) pointed out that at temperatures below 20 K, some of the accreted carbon atoms are quickly converted to formaldehyde. The remaining physisorbed atoms are partially changed to chemisorbed and hydrogenated to CH4. Physisorbed carbon atoms can diffuse at temperatures between 20 K and 30 K, promoting insertion reactions with other carbon atoms or species. Future observations with the James Webb Space Telescope (JWST) could provide additional details about the significance of chemisorption in forming COMs in the ice phase.
4.2 Implication of the computed BEs on Astrochemical model
Recently, using ALMA observations, Coutens et al. (2019) detected multiple transitions of HONO in the low-mass protostellar system IRAS 16293–2422. This system consists of three protostars, two located toward source A and one toward source B (Maureira et al., 2020). The identification by Coutens et al. (2019) specifically focuses on the protostar associated with source B. They also attempted to explain the abundances of HONO-related species and their ratios using a chemical model. However, they were unable to account for all the observed features. In their study, Coutens et al. (2019) estimated abundance ratios of HONO/NH2OH and HONO/HNO to be greater than or equal to and , respectively. Using their chemical model, they successfully explained the abundance of HONO, N2O, and NO2. However, their model resulted in an overproduction of NH2OH and HNO, leading to lower HONO/NH2OH () and HONO/HNO (0.01) ratios compared to the observed values. Our study suggests a weak binding of NO with the water substrate. It would be interesting to see how the estimated BEs would affect the modeled abundances of these species.
We utilize the Rokko code Furuya et al. (2015, 2017) for our chemical modeling, which encompasses three phases: gas, icy grain surface, and homogeneous bulk ice mantle. Our active ice layers are limited to the top four monolayers. All layers, including the active ice layers, are considered part of the bulk ice mantle. The gas and ice phase networks are based on Garrod (2013). The active layers undergo photodissociation and photodesorption of species. A constant reactive desorption of all the species is assumed with a factor of 0.01, resulting in a single product. In molecular dynamics studies, it had been determined that approximately 10% of the formation energy is retained after the formation of H2S from HS+H (Bariosco et al., 2024, and references therein). However, based on experimentally determined reaction desorption probabilities for H2S, HDS, and D2S (Oba et al., 2019; Furuya et al., 2022), a reactive desorption factor of 0.03 is being utilized. A standard cosmic-ray ionization rate of s-1 is considered.
4.2.1 Dark Cloud
For the dark cloud model, we start with a hydrogen number density, cm-3, visual extinction () of 10 mag, and a gas and ice temperature of 10 K. The initial elemental abundances of species are taken from Garrod et al. (2017). The upper panel of Fig. 5 shows the time evolution of the radicals. Here, “G” refers to the notation when the whole BE sets are from Garrod (2013), whereas “G+S” refers to the BEs for the radicals taken from this work, and the rest are considered from Garrod (2013). In this work, we are using a default diffusion-to-BE ratio of 0.4. Since no estimated BEs were available for HONO, H2NO2, and H3NO2, we used our estimated values of 2918 K, 4205 K, and 3044 K, respectively, in both “G” and “G+S.” It shows that the changes in the BE would have a minor effect on the gas phase abundances of these radicals. However, it shows a major effect on the ice phase abundances of these radicals, especially for the ice phase NO and CH. NO ice abundance is almost reduced by a factor of at years. Meanwhile, the ice phase abundance of CH increased by a factor of 100. The reduction in the abundance of NO is due to the weak binding of NO obtained with this work ( K) compared to the earlier estimate (1600 K; Wakelam et al., 2017). The increase in the surface abundance of CH radicals is due to the estimate of a higher BE ( K) than previous work (Garrod, 2013, 925 K). The shaded lines on the graph indicate the uncertainty of our BE values from Table 2. Since NO has a very weak binding with the water substrate, its small uncertainty (about 94 K) in the BE estimation could cause a significant change in its ice phase abundance (see the black shaded area in the top right panel of Fig. 5).
The lower panel of Fig. 5 shows the time evolution of the NO-related species. It clearly shows that with the present BE estimation of NO, we could expect a favorable formation of gas phase HONO in a dark cloud. On the contrary, the formation of gas phase NH2OH is heavily affected, which could be useful in explaining the observed HONO/NH2OH ratio in star-forming regions (Coutens et al., 2019) discussed in 4.2.2. The shaded regions represent when the uncertainty in the BE calculations is considered. The obtained abundances and abundance ratio with HONO are noted in Table 4. Due to the weak binding of NO obtained in this study, we have a higher abundance of the ice phase NO2 through the reaction: , resulting in a favorable formation of HONO-gr by . The formation of NH2OH-gr process through successive hydrogenation reactions: . When our BE value of NO is used, due to the high abundance of atomic oxygen on the grain compared to the atomic hydrogen, NO-gr is mainly channelized to form , resulting in a decrease in the formation of . On the contrary, it facilitates the HONO-gr formation. In the lower panel of Fig. 5, we have observed significant uncertainties in the ice phase abundances of HNO and NH2OH due to the uncertainty surrounding the BE of NO ( 94 K).
Species | Dark cloud | Hot core | Observation |
HONO | — | ||
NO | — | ||
N2O | — | ||
HNO | — | ||
NO2 | — | ||
NH2OH | — | ||
CH3OH | — | ||
HONO/NO | |||
HONO/N2O | |||
HONO/HNO | |||
HONO/NO2 | |||
HONO/NH2OH | |||
HONO/CH3OH |
g when the BEs from Garrod (2013) are used.
s when the BEs calculated here are used and for others Garrod (2013) used.
4.2.2 Hot corino
We have further examined a low-mass star-forming region to assess the impact of our proposed BEs on abundance estimation. This study considers three stages: the isothermal collapsing stage, the warm-up stage, and the post-warm-up stage. During the isothermal collapsing stage, the cloud is allowed to collapse from a minimum density of cm-3 to a maximum density of cm-3 in one million years, while maintaining a gas temperature of 10 K and ice temperature of 10 K. The initial abundance for this model is taken from the final abundance of the dark cloud model discussed in 4.2.1. The visual extinction at the first stage can vary with the density following the relation provided by Garrod & Pauly (2011), resulting in a variation of . In the warm-up stage, the density and visual extinction parameters are consistently set to their maximum values for an additional one million years while the temperature gradually rises to a peak of K. During the post-warm-up stage, the parameters listed above remain at their maximum values for an additional years.
The upper panel of Fig. 6 shows the time evolution of the abundances of gas-phase radicals. With the new BE value, the gas-phase abundance of NO (black dashed curve) has decreased significantly during the warm-up stage. In the lower panel of Fig. 6, we show the abundances of NO and its related species separately. The uncertainties in the BE calculations are also considered here as it was discussed in the context of Fig. 5. In Table 4, we compare the peak abundances (peak value is taken beyond the isothermal stage) and the peak abundance ratio of some NO-related species obtained from our simulations with that observed by Coutens et al. (2019). With the new BEs (represented by the dashed curves), we observe a significant reduction in the abundances of NO, N2O, HNO, and NH2OH. Conversely, there is a surge in the abundance of HONO.
NH2OH is primarily produced in the cold ice phase through the successive sequential hydrogenation of NO-gr, as discussed in 4.2.1, and it can sublimate to the gas phase after around years, depending on its BE considered (6806 K; Garrod, 2013). Due to the weak binding of NO, its abundance in the ice phase decreases as it forms NO2, leading to a decrease in NH2OH. We estimated a BE of 2918 K for HONO, so it is expected to be released earlier than NH2OH in the gas phase. However, an additional radical-radical pathway () is responsible for ice phase HONO production in warmer regions. The low BE of NO obtained here further facilitates this radical-radical ice phase reaction.
In our study, with our estimated BE of NO, we achieve a lower abundance of NH2OH and HNO (see Fig. 6 and Table 4). We obtain peak abundance ratios of HONO/NH2OH and HONO/HNO of 143 and 3781, respectively, consistent with the observation. Table 4 depicts that the other ratios from the observations seem to be between the previous model (Coutens et al., 2019) and this model. This may indicate that the diffusion activation energy of NO was overestimated in the previous one and underestimated in the present.
5 Conclusions
This study aims to calculate the BE of certain interstellar radicals on cluster surfaces of ASW. The work employs a fully quantum chemical approach to examine a large number of binding sites for all the considered radicals using a suitable ASW cluster model. The significant findings of our work are presented below:
1.
We estimate the binding energies (BEs) of 7 diatomic radicals using a pure quantum chemical approach with high-level water clusters and methods, and discuss the implications of the computed BEs on astrochemical models.
2.
In our study, the BSSE correction is carefully employed extensively using the CP method. This correction is crucial, especially with small and medium-sized basis sets, for obtaining accurate and reliable estimates of the BEs between radical species and the ASW cluster.
3.
Our results suggest that most of the time, we have physisorption, indicating weak bonding between radicals and the water molecule. However, in addition to physisorption, we find chemisorption of CH with cluster in two cases. In this case, a chemical bond is formed between CH and an oxygen atom of a water molecule. Figure 9 illustrates the production of CH2OH through the chemisorption of CH on a water substrate. This warrants the potential formation of methanol through hydrogenation in warmer regions. Future observations with the JWST could address this issue. Moreover, our estimated value of the BE (physisorption) of CH is 2-3 times higher than the available guess value. Hence, we suggest using these values for astrochemical models.
4.
We obtain a significantly lower BE value of NO (704 K), which is in excellent agreement with a very recent study (average value of 712 K with ASW model; Martínez-Bachs et al., 2024) compared to that used earlier (1600 K). Our obtained BEs of NO would be useful in explaining the observed HONO/NH2OH and HONO/HNO ratio obtained in the B component of the low-mass protostellar binary IRAS 16293–2422.
Acknowledgements.
The authors thank the referee for the valuable insights and constructive comments that significantly improved the manuscript. This work is completed while M.S. is a postdoctoral fellow at the Institute of Planetology and Astrophysics of Grenoble (IPAG), University of Grenoble Alpes. M.S. acknowledges financial support through the European Research Council (consolidated grant COLLEXISM, grant agreement ID: 811363). N.I. acknowledges the support of FONDECYT Regular Grant 1241193 and Vicerectoría de Investigación y Postgrado (VRIP). A.D. acknowledges the support of MPE for sponsoring a scientific visit to MPE.References
- Aikawa et al. (1996) Aikawa, Y., Miyama, S. M., Nakano, T., & Umebayashi, T. 1996, ApJ, 467, 684
- Bariosco et al. (2024) Bariosco, V., Pantaleone, S., Ceccarelli, C., et al. 2024, Monthly Notices of the Royal Astronomical Society, 531, 1371
- Boogert et al. (2015) Boogert, A. C. A., Gerakines, P. A., & Whittet, D. C. B. 2015, ARA&A, 53, 541
- Bovolenta et al. (2020) Bovolenta, G., Bovino, S., Vöhringer-Martinez, E., et al. 2020, Molecular Astrophysics, 21, 100095
- Bovolenta et al. (2022) Bovolenta, G. M., Vogt-Geisse, S., Bovino, S., & Grassi, T. 2022, The Astrophysical Journal Supplement Series, 262, 17
- Boys & Bernardi (1970) Boys, S. & Bernardi, F. 1970, Molecular Physics, 19, 553
- Chai & Head-Gordon (2008) Chai, J.-D. & Head-Gordon, M. 2008, Physical Chemistry Chemical Physics (Incorporating Faraday Transactions), 10, 6615
- Coutens et al. (2019) Coutens, A., Ligterink, N. F. W., Loison, J.-C., et al. 2019, A&A, 623, L13
- Das et al. (2008a) Das, A., Acharyya, K., Chakrabarti, S., & Chakrabarti, S. K. 2008a, A&A, 486, 209
- Das et al. (2010) Das, A., Acharyya, K., & Chakrabarti, S. K. 2010, MNRAS, 409, 789
- Das & Chakrabarti (2011) Das, A. & Chakrabarti, S. K. 2011, MNRAS, 418, 545
- Das et al. (2019) Das, A., Gorai, P., & Chakrabarti, S. K. 2019, A&A, 628, A73
- Das et al. (2015) Das, A., Majumdar, L., Sahu, D., et al. 2015, ApJ, 808, 21
- Das et al. (2016) Das, A., Sahu, D., Majumdar, L., & Chakrabarti, S. K. 2016, MNRAS, 455, 540
- Das et al. (2018) Das, A., Sil, M., Gorai, P., Chakrabarti, S. K., & Loison, J. C. 2018, ApJS, 237, 9
- Duflot et al. (2021) Duflot, D., Toubin, C., & Monnerville, M. 2021, Frontiers in Astronomy and Space Sciences, 8, 24
- Dulieu et al. (2013) Dulieu, F., Congiu, E., Noble, J., et al. 2013, Scientific Reports, 3, 1338
- Dunning (2000) Dunning, T. H. 2000, The Journal of Physical Chemistry A, 104, 9062
- Enrique-Romero et al. (2022) Enrique-Romero, J., Rimola, A., Ceccarelli, C., et al. 2022, ApJS, 259, 39
- Ferrero et al. (2020) Ferrero, S., Zamirri, L., Ceccarelli, C., et al. 2020, The Astrophysical Journal, 904, 11
- Frisch et al. (2009) Frisch, M. J., Trucks, G. W., Schlegel, H. B., et al. 2009, Gaussian 09 Revision A.01, gaussian Inc. Wallingford CT
- Frisch et al. (2013) Frisch, M. J., Trucks, G. W., Schlegel, H. B., et al. 2013, Gaussian 09 Revision D.01, gaussian Inc. Wallingford CT
- Frisch et al. (2016) Frisch, M. J., Trucks, G. W., Schlegel, H. B., et al. 2016, Gaussian 16 Revision B.01, gaussian Inc. Wallingford CT
- Frisch et al. (2019) Frisch, M. J., Trucks, G. W., Schlegel, H. B., et al. 2019, Gaussian 16 Revision C.01, gaussian Inc. Wallingford CT
- Furuya et al. (2015) Furuya, K., Aikawa, Y., Hincelin, U., et al. 2015, Astronomy and Astrophysics, 584, A124
- Furuya et al. (2017) Furuya, K., Drozdovskaya, M. N., Visser, R., et al. 2017, Astronomy and Astrophysics, 599, A40
- Furuya et al. (2022) Furuya, K., Oba, Y., & Shimonishi, T. 2022, The Astrophysical Journal, 926, 171
- Garrod (2013) Garrod, R. T. 2013, ApJ, 765, 60
- Garrod et al. (2017) Garrod, R. T., Belloche, A., Müller, H. S. P., & Menten, K. M. 2017, A&A, 601, A48
- Garrod & Herbst (2006) Garrod, R. T. & Herbst, E. 2006, A&A, 457, 927
- Garrod & Pauly (2011) Garrod, R. T. & Pauly, T. 2011, ApJ, 735, 15
- Garrod et al. (2008) Garrod, R. T., Widicus Weaver, S. L., & Herbst, E. 2008, ApJ, 682, 283
- Germain et al. (2022) Germain, A., Tinacci, L., Pantaleone, S., Ceccarelli, C., & Ugliengo, P. 2022, ACS Earth and Space Chemistry, 6, 1286
- Ghosh et al. (2022) Ghosh, R., Sil, M., Kumar Mondal, S., et al. 2022, Research in Astronomy and Astrophysics, 22, 065021
- Gibb et al. (2004) Gibb, E. L., Whittet, D. C. B., Boogert, A. C. A., & Tielens, A. G. G. M. 2004, ApJS, 151, 35
- Grieco et al. (2024) Grieco, F., Dulieu, F., De Looze, I., & Baouche, S. 2024, MNRAS, 527, 10604
- Grieco et al. (2023) Grieco, F., Theulé, P., De Looze, I., & Dulieu, F. 2023, Nature Astronomy, 7, 541
- Hama & Watanabe (2013) Hama, T. & Watanabe, N. 2013, Chemical Reviews, 113, 8783
- Hasegawa & Herbst (1993) Hasegawa, T. I. & Herbst, E. 1993, MNRAS, 261, 83
- Hasegawa et al. (1992) Hasegawa, T. I., Herbst, E., & Leung, C. M. 1992, ApJS, 82, 167
- He et al. (2016) He, J., Acharyya, K., & Vidali, G. 2016, ApJ, 825, 89
- He & Vidali (2014) He, J. & Vidali, G. 2014, ApJ, 788, 50
- Hendrix et al. (2023) Hendrix, J., Bera, P. P., Lee, T. J., & Head-Gordon, M. 2023, Molecular Physics, 0, e2252100
- Ioppolo et al. (2021) Ioppolo, S., Fedoseev, G., Chuang, K. J., et al. 2021, Nature Astronomy, 5, 197
- Jensen (2024) Jensen, F. 2024, Journal of Chemical Theory and Computation, 20, 767, pMID: 38174405
- Jin & Garrod (2020) Jin, M. & Garrod, R. T. 2020, The Astrophysical Journal Supplement Series, 249, 26
- Keane et al. (2001) Keane, J. V., Tielens, A. G. G. M., Boogert, A. C. A., Schutte, W. A., & Whittet, D. C. B. 2001, A&A, 376, 254
- Martínez-Bachs et al. (2024) Martínez-Bachs, B., Ferrero, S., Ceccarelli, C., Ugliengo, P., & Rimola, A. 2024, ApJ, 969, 63
- Martínez-Bachs et al. (2020) Martínez-Bachs, B., Ferrero, S., & Rimola, A. 2020, in Computational Science and Its Applications – ICCSA 2020, ed. O. Gervasi, B. Murgante, S. Misra, C. Garau, I. Blečić, D. Taniar, B. O. Apduhan, A. M. A. C. Rocha, E. Tarantino, C. M. Torre, & Y. Karaca (Cham: Springer International Publishing), 683–692
- Maureira et al. (2020) Maureira, M. J., Pineda, J. E., Segura-Cox, D. M., et al. 2020, ApJ, 897, 59
- McGuire (2022) McGuire, B. A. 2022, ApJS, 259, 30
- Minissale et al. (2022) Minissale, M., Aikawa, Y., Bergin, E., et al. 2022, ACS Earth and Space Chemistry, 6, 597
- Miyazaki et al. (2020) Miyazaki, A., Watanabe, N., Sameera, W. M. C., et al. 2020, Phys. Rev. A, 102, 052822
- Molpeceres et al. (2021) Molpeceres, G., Kästner, J., Fedoseev, G., et al. 2021, The Journal of Physical Chemistry Letters, 12, 10854, pMID: 34727500
- Mondal et al. (2021) Mondal, S. K., Gorai, P., Sil, M., et al. 2021, The Astrophysical Journal, 922, 194
- Nguyen & Peeters (2021) Nguyen, T. L. & Peeters, J. 2021, Phys. Chem. Chem. Phys., 23, 16142
- Oba et al. (2019) Oba, Y., Tomaru, T., Kouchi, A., & Watanabe, N. 2019, The Astrophysical Journal, 874, 124
- Penteado et al. (2017) Penteado, E. M., Walsh, C., & Cuppen, H. M. 2017, ApJ, 844, 71
- Perrero et al. (2022) Perrero, J., Enrique-Romero, J., Ferrero, S., et al. 2022, The Astrophysical Journal, 938, 158
- Phelps & Dalby (1966) Phelps, D. H. & Dalby, F. W. 1966, Phys. Rev. Lett., 16, 3
- Piacentino & Öberg (2022) Piacentino, E. L. & Öberg, K. I. 2022, ApJ, 939, 93
- Rimola et al. (2018) Rimola, A., Skouteris, D., Balucani, N., et al. 2018, ACS Earth and Space Chemistry, 2, 720
- Sameera et al. (2021) Sameera, W. M. C., Senevirathne, B., Andersson, S., et al. 2021, Journal of Physical Chemistry A, 125, 387
- Sameera et al. (2017) Sameera, W. M. C., Senevirathne, B., Andersson, S., Maseras, F., & Nyman, G. 2017, The Journal of Physical Chemistry C, 121, 15223
- Sheng et al. (2011) Sheng, X. W., Mentel, L., Gritsenko, O. V., & Baerends, E. J. 2011, Journal of Computational Chemistry, 32, 2896
- Shimonishi et al. (2018) Shimonishi, T., Nakatani, N., Furuya, K., & Hama, T. 2018, ApJ, 855, 27
- Shingledecker et al. (2019) Shingledecker, C. N., Vasyunin, A., Herbst, E., & Caselli, P. 2019, The Astrophysical Journal, 876, 140
- Sil et al. (2018) Sil, M., Gorai, P., Das, A., et al. 2018, ApJ, 853, 139
- Sil et al. (2017) Sil, M., Gorai, P., Das, A., Sahu, D., & Chakrabarti, S. K. 2017, European Physical Journal D, 71, 45
- Sil et al. (2021) Sil, M., Srivastav, S., Bhat, B., et al. 2021, AJ, 162, 119
- Srivastav et al. (2022) Srivastav, S., Sil, M., Gorai, P., et al. 2022, Monthly Notices of the Royal Astronomical Society, 515, 3524
- Tinacci et al. (2023) Tinacci, L., Germain, A., Pantaleone, S., et al. 2023, The Astrophysical Journal, 951, 32
- Tinacci et al. (2022) Tinacci, L., Germain, A., Pantaleone, S., et al. 2022, ACS Earth and Space Chemistry, 6, 1514
- Tsuge et al. (2023) Tsuge, M., Molpeceres, G., Aikawa, Y., & Watanabe, N. 2023, Nature Astronomy, 7, 1351–1358
- Wakelam et al. (2017) Wakelam, V., Loison, J. C., Mereau, R., & Ruaud, M. 2017, Molecular Astrophysics, 6, 22
- Whittet (2003) Whittet, D. 2003, Dust in the galactic environment, by DCB Whittet
- Zhurko (2005) Zhurko, G. A. 2005, ivanovo, Russia
Appendix A Optimized geometries and calculated BE values
Geometry optimization of the ASW clusters and the BE values estimation of the 7 radicals are performed using DFT-B97X-D/6-311+G(d,p) level theory as stated in Section 2.
ASW | -H binding | Optimized energy⋆ | Relative energy | Optimized cluster |
clusters | sites avaiable | (Hartrees) | [Eref.= ASW(1)] (Kelvin) | structures |
ASW(1) | 7 | -1528.510023 | 0 | |
ASW(2) | 7 | -1528.499694 | 3262 | |
ASW(3) | 9 | -1528.505233 | 1513 | |
ASW(4) | 10 | -1528.494038 | 5048 | |
ASW(5) | 8 | -1528.499406 | 3353 | |
ASW(6) | 9 | -1528.494265 | 4976 | |
ASW(7) | 8 | -1528.508844 | 372 | |
ASW(8) | 9 | -1528.503436 | 2080 | |
ASW(9) | 8 | -1528.500329 | 3061 |
⋆ Optimized energy values correspond to the ZPE-corrected electronic energy in Hatree.
The sign indicates the stability of the structure; the larger it is, the more stable the optimized geometry.
-H binding sites | BE values (Kelvin) | Highest BE value (Kelvin) | ||
BSSE | BSSE | BSSE | BSSE | |
ASW(1) | ||||
1 | 2465 | 2162 | ||
2 | 2955 | 2674 | ||
3 | 1880 | 1500 | 3990 | 3643 |
4 | 3125 | 2816 | ||
5 | 3990 | 3643 | (S.F. ) | |
6 | 3116 | 2793 | ||
7 | 1662 | 1452 | ||
ASW(2) | ||||
1 | 1709 | 1495 | ||
2 | 1498 | 1265 | ||
3 | 2887 | 2553 | ||
4 | 4734 | 4296 | 4734 | 4296 |
5 | 2957 | 2630 | ||
6 | 3456 | 3130 | ||
7 | 4349 | 3914 | ||
ASW(3) | ||||
1 | 3410 | 3063 | ||
2 | 3435 | 2931 | ||
3 | 2939 | 2640 | ||
4 | 1683 | 1456 | ||
5 | 5580 | 5028 | 5580 | 5028 |
6 | 3906 | 3482 | ||
7 | 1758 | 1532 | ||
8 | 3346 | 2999 | ||
9 | 4492 | 4080 | ||
ASW(4) | ||||
1 | 3676 | 3331 | ||
2 | 3948 | 3481 | ||
3 | 3219 | 2888 | ||
4 | 2964 | 2508 | ||
5 | 1799 | 1573 | 4719 | 4261 |
6 | 4114 | 3654 | ||
7 | 3170 | 2867 | ||
8 | 4062 | 3579 | ||
9 | 4719 | 4261 | ||
10 | 3012 | 2687 | ||
ASW(5) | ||||
1 | 3582 | 3246 | ||
2 | 3909 | 3572 | ||
3 | 3150 | 2801 | ||
4 | 3642 | 3294 | ||
5 | 3561 | 3237 | 5397 | 4902 |
6 | 1914 | 1683 | ||
7 | 4285 | 3882 | ||
8 | 5397 | 4902 | ||
ASW(6) | ||||
1 | 2785 | 2320 | ||
2 | 3497 | 3168 | ||
3 | 3629 | 3300 | ||
4 | 4178 | 3774 | ||
5 | 4770 | 4191 | 4802 | 4366 |
6 | 4569 | 4046 | ||
7 | 3755 | 3399 | ||
8 | 4802 | 4366 | ||
9 | 4184 | 3780 | ||
ASW(7) | ||||
1 | 3798 | 3337 | ||
2 | 3468 | 3134 | ||
3 | 3525 | 3190 | ||
4 | 3340 | 3005 | ||
5 | 2605 | 2104 | 3798 | 3337 |
6 | 3560 | 3229 | ||
7 | 3315 | 2967 | ||
8 | 2972 | 2533 | ||
ASW(8) | ||||
1 | 4436 | 3986 | ||
2 | 1633 | 1428 | ||
3 | 2990 | 2586 | ||
4 | 3609 | 3270 | ||
5 | 3237 | 2875 | 4436 | 3986 |
6 | 3008 | 2686 | ||
7 | 3414 | 3077 | ||
8 | 2923 | 2618 | ||
9 | 3642 | 3226 | ||
ASW(9) | ||||
1 | 2803 | 2492 | ||
2 | 3236 | 2931 | ||
3 | 4972 | 4503 | ||
4 | 3132 | 2807 | ||
5 | 5216 | 4761 | 5216 | 4761 |
6 | 2751 | 2329 | ||
7 | 3560 | 3225 | ||
8 | 1746 | 1518 | ||
Total 75 sites | — | — | Average |
-H binding | BE values (Kelvin) | Highest BE value (Kelvin) | Scaled by | |||
sites of ASW(1) | BSSE | BSSE | BSSE | BSSE | 1.188 for BSSE | 1.177 for BSSE |
CH radical | ||||||
1 | 1837 | 1670 | ||||
2 | 1688 | 1530 | ||||
3 | 1610 | 1454 | ||||
4 | 1601 | 1446 | 1892 | 1737 | 2248 | 2044 |
5 | 31315∗ | — | ||||
6 | 1892 | 1737 | ||||
7 | 14746∗ | — | ||||
NH radical | ||||||
1 | 1441 | 1282 | ||||
2 | 1409 | 1258 | ||||
3 | 1363 | 1223 | ||||
4 | 1331 | 1182 | 2083 | 1898 | 2475 | 2234 |
5 | 2083 | 1898 | ||||
6 | 1569 | 1421 | ||||
7 | 1425 | 1280 | ||||
SH radical | ||||||
1 | 1661 | 1363 | ||||
2 | 1870 | 1610 | ||||
3 | 1010 | 866 | ||||
4 | 1839 | 1572 | 2416 | 2109 | 2870 | 2482 |
5 | 2416 | 2109 | ||||
6 | 1100 | 942 | ||||
7 | 1276 | 1131 | ||||
CN radical | ||||||
1 | 1180 | 1068 | ||||
2 | 1192 | 1086 | ||||
3 | 1121 | 1016 | ||||
4 | 1028 | 925 | 1192 | 1086 | 1416 | 1278 |
5 | 6652∗ | 6317∗ | ||||
6 | 1152 | 1047 | ||||
7 | 1146 | 1044 | ||||
NS radical | ||||||
1 | 2117 | 1735 | ||||
2 | 2038 | 1677 | ||||
3 | 1808 | 1473 | ||||
4 | 2316 | 1879 | 2606 | 2225 | 3096 | 2619 |
5 | 2606 | 2225 | ||||
6 | 1865 | 1513 | ||||
7 | 2462 | 2039 | ||||
NO radical | ||||||
1 | 686 | 510 | ||||
2 | 540 | 353 | ||||
3 | 895 | 591 | ||||
4 | 931 | 598 | 931 | 598 | 1106 | 704 |
5 | 743 | 574 | ||||
6 | 658 | 490 | ||||
7 | 384 | 259 |
-H | B97X-D / 6-311+G(d,p) | ||
binding sites | BSSE | BSSE | |
1 | 2465 | 2162 | 1.1401 |
2 | 2955 | 2674 | 1.1051 |
3 | 1880 | 1500 | 1.1253 |
4 | 3125 | 2816 | 1.1097 |
5 | 3990 | 3643 | 1.0952 |
6 | 3116 | 2793 | 1.1156 |
7 | 1662 | 1452 | 1.1446 |